OpenStudy (anonymous):

(3,4) and (9,6) answer in slope intercept from y=mx+b

OpenStudy (anonymous):

slope = 1/3 y - y1 = m (x - x1) y - 4 = (1/3)(x-3) y = (x/3) - 1 + 4 y = (x/3) + 3